OpenStudy (anonymous):

a rectangular dance floor measures 24ft by 32ft . The length and width of the floor will both be increased by x feet. write an equation that can be used to determine the value of x, in feet, if the area of new dance floor is 1,174.25 square feet

OpenStudy (amistre64):

24x*32x = 1174.25 768x^2 = 1174.25

OpenStudy (anonymous):

what are the dimensions of the new dance floor ,in feet, if the area is 1,174.25 square feet

OpenStudy (amistre64):

x = sqrt(1174.25/768) sooo 24+x by 32+x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

yw :)